### v1

@@ 4.6 Ablation Study
Table 3: Performance evaluation of QLFR-CML and its variant on short text datasets. The best results are in bold. ‘w/o ECCA’ illustrates results without employing the ECCA strategy, while ‘w/o SSE-CoT’ and ‘w/o DA-CoT’ display outcomes when omitting SSE-CoT and DA-CoT rationales, respectively.
QLFR-CML Ablation. Ablation studies were performed on six benchmark datasets to assess the influence of particular strategies or components on the QLFR-CML method. We developed three variations of the QLFR-CML, as follows:
  • QLFR-CML+w/o ECCA: During the fine-tuning stage, the original text $X$ is used as input to predict the label without employing the ECCA strategy.
  • QLFR-CML+w/o SSE-CoT: The task of generate SSE-CoT rationales is removed from the multi-task learning process.
  • QLFR-CML+w/o DA-CoT: Similarly, the task of generate DA-CoT rationales is excluded from the multi-task learning process.
Table 3 displays the outcomes of our ablation study, highlighting the best-performing metrics in bold. These findings demonstrate the integral role of each strategy and component in the effectiveness of our QLFR-CML method. Notably, ECCA contributes most substantially, with SSE-CoT and DA-CoT also providing significant enhancements. The significant improvement provided by ECCA confirms the importance of a prompt for generative models with an encoder-decoder architecture. The improvement observed with SSE-CoT and DA-CoT affirms that our strategy effectively transfers the knowledge and capabilities of LLM to a smaller model.
@@
Table 4: Performance evaluation of QLFR and its variant on short text datasets. The best results are in bold. ‘w/o rewriting’ and ‘w/o retrieval’ refer to results without the rewriting step and retrieval step in SSE-CoT, respectively. ‘ w/o both’ refers to the exclusion of both steps
QLFQ Ablation. Ablation studies were performed on six benchmark datasets to assess the influence of the components of the QLFQ framework. We developed two variations of the QLFQ as follows.
  • QLFR+w/o rewriting: Omission of the rewriting step, using the concatenation of the original input with the retrieved text as input.
  • QLFR+w/o retrieval: Omission of the retrieval step, employing rewritten original inputs as input.
  • QLFR+w/o both: Utilizing the original input without modification.
@@
4.7 Analysis of In-Context Learning for QLFR
In addition to the supervised fine-tuning (SFT) paradigm, in-context learning has gained popularity. We chose four LLMs, ChatGLM, LLaMA2-7B, LLaMA2-13B, and GPT-3, for experimentation in the zero-shot and one-shot settings. Given the rigorous demands of in-context learning, the QLFR was tested solely on the GPT-3. We manually selected a sample from each category in a one-shot setting and constructed the SSE-CoT as the context. Specific examples of the input formats are provided in Appendix B. The experimental results are presented in Figure 5.
@@ 4.7 Analysis of In-Context Learning for QLFR
Zero-shot setting. In the analysis across the three datasets under a zero-shot setting, the in-context learning abilities of the four models were ranked as follows: GPT-3 outperformed LLaMA2-13B, which, in turn, surpassed both ChatGLM and LLaMA2-7B. This hierarchy is attributable to in-context learning aptitude inherent in larger models, with increased parameters correlating with enhanced performance. Applying our QLFR method to GPT-3 resulted in performance gains across the board. For instance, the ACC on the Ohsumed dataset improved from 51.6% to 52.9%. These findings indicate that our approach is not limited to the SFT but is also applicable to the context learning paradigm.
@@
Comparing the results of the zero-shot and one-shot tasks, contrary to our initial expectations, the hypothesis that providing task-relevant examples enhances the comprehension and task response of LLM is not supported. The findings reveal that only the GPT improves performance in one-shot learning compared with zero-shot learning in the two datasets. In contrast, a decline is observed in the remaining datasets. Further analysis of the model outputs suggested that LLM with approximately 10 billion parameters tended to demonstrate a reduced capacity for processing instructions with greater input complexity. In contrast, the GPT consistently exhibits a robust understanding of instructions and sustains its performance despite increased context length.
Comparison of the results of SFT and in-context learning paradigm. GPT-3, with over a hundred billion parameters, exhibits a remarkable in-context learning performance. In the one-shot setting, GPT-3 exceeded QLFR-CML, and GPT-3+QLFR nearly matched the supervised QLFR method. However, for models with billions of parameters, the SFT paradigm remained predominant.
4.8 Analysis of different base models for QLFR-CML
In our QLFR-CML method, the initial stage involves the extraction of rationales with the assistance of an LLM, followed by a fine-tuning phase that requires a smaller model. Therefore, we investigated the outcomes of employing the same LLM with various smaller models and the results of utilizing the same smaller model with different LLMs.
@@ 4.8 Analysis of different base models for QLFR-CML
Different smaller models. As shown in Figure 6, we utilized LLaMA2-13B as the LLM and three versions of Flan-T5-Base, Flan-T5-Large, and Flan-T5-XL as smaller models, with parameter counts of 250M, 780M, and 3 B, respectively. It is apparent that, with an increase in the parameter size of the smaller models, there is a corresponding enhancement in both the ACC and F1 metrics, which suggests a correlation between the capacity of the smaller model and its performance. However, this relationship is not strictly linear. For instance, in the Ohsumed dataset, the increase in ACC from the 250M model to the 780M model does not reflect the expected proportional improvement, indicating diminishing returns as the model size increases.
@@
4.9 Analysis of different prompts for QLFR-CML
@@
Table 6: Evaluation of Different Prompts in QLFR-CML: the efficacy of ECCA demonstrated through significant outcomes.
Table 6 lists several interesting trends. Prompts enhanced performance across all measured metrics relative to their absence. Furthermore, our method outperformed the less semantically rich prompt1 and demonstrated greater accuracy than prompt2 in the two datasets. Although our approach does not outperform prompt2 on the MR dataset in terms of ACC, it offers considerable benefits. It obviates the necessity for creating manual, dataset-specific prompts, which yield notably shorter input lengths than prompt2 by simply appending label names, benefiting processing efficiency and model scalability.
@@ 4.10 Analysis of training size
The results presented in Figure 7 show that the proposed method significantly outperforms the other two models in low-resource environments. The data revealed a consistent increase in test accuracy correlated with the rising proportion of training data for both the MR and TagMyNews datasets, consistent with supervised learning principles. Additionally, the acceleration or plateau of the ACC improvement suggests diminishing returns in the model performance beyond a certain data threshold. Notably, the ACC of SHINE and ChatGLM approach parity, whereas our methods consistently outperform them, indicating that, with adequate data, the inherent strengths of the model become increasingly influential.
@@ 5 Conclusion
In this study, we introduce a novel reasoning framework, Quartet Logic: a Four-Step Reasoning (QLFR) framework, alongside the CoT-Driven Multi-task learning (QLFR-CML) method, to address the challenge of short text classification. The QLFR framework employs large Language Models (LLMs) for resolution, whereas the QLFR-CML method uses smaller models. Our framework leverages extensive, inherent knowledge and diverse abilities that surpass prior GCN-based models. To harness this knowledge within LLMs, we implemented syntactic and semantic enrichment CoTs to facilitate methodological reasoning by LLMs. Moreover, to impart this knowledge to the smaller models, we adopted a multi-task learning approach. We introduced an additional Domain-Adaptive CoT to enrich the knowledge transfer. In the fine-tuning phase for smaller models, we proposed an Explicit Category Context Augmentation mechanism to improve task comprehension. Comprehensive experiments were conducted across six prevalent datasets to determine the effectiveness of the proposed framework. The experimental results demonstrated that the proposed framework significantly outperformed established baselines. However, this complex task remains unresolved because its performance on domain-specific datasets remains suboptimal. Future efforts should focus on integrating LLMs and additional knowledge to refine the proposed approach.
@@ 6 Limitations
Our study involved carefully designed experiments to validate the effectiveness of the proposed method. However, limited hardware resources and time constraints affect the ideal number of LLMs employed. Moreover, swift progress in large-language-model technologies poses the challenge of maintaining up-to-date comparative analyses. Nonetheless, we endeavored to ensure the robustness and relevance of our findings within these limitations.
@@ Appendix B In-Context Learning input apllications
Appendix C Different Prompts for QLFR-CML
Categorize this text: ‘wal-mart buys social media firm kosmix’.
Given the short text ‘wal-mart buys social media firm kosmix’, classify it into one of the categories. The categories are health, sport, entertainment, business, sci_tech, U.S. and world.

### v2

@@ 4.6 Ablation Study
Table 3: Performance evaluation of CDMT and its variant on short text datasets. The best results are in bold. ‘w/o ECCA’ illustrates results without employing the ECCA strategy, while ‘w/o SSE-CoT’ and ‘w/o DA-CoT’ display outcomes when omitting SSE-CoT and DA-CoT rationales, respectively.
CDMT Ablation. Ablation studies were performed on six benchmark datasets to assess the influence of particular strategies or components on the CDMT method. We developed three variations of the CDMT, as follows:
  • CDMT+w/o ECCA: During the fine-tuning stage, the original text $X$ is used as input to predict the label without employing the ECCA strategy.
  • CDMT+w/o SSE-CoT: The task of generate SSE-CoT rationales is removed from the multi-task learning process.
  • CDMT+w/o DA-CoT: Similarly, the task of generate DA-CoT rationales is excluded from the multi-task learning process.
Table 3 displays the outcomes of our ablation study, highlighting the best-performing metrics in bold. These findings demonstrate the integral role of each strategy and component in the effectiveness of our CDMT method. Notably, ECCA contributes most substantially, with SSE-CoT and DA-CoT also providing significant enhancements. The significant improvement provided by ECCA confirms the importance of a prompt for generative models with an encoder-decoder architecture. The improvement observed with SSE-CoT and DA-CoT affirms that our strategy effectively transfers the knowledge and capabilities of LLM to a smaller model.
@@
Table 4: Performance evaluation of SSE-CoT and its variant on short text datasets. The best results are in bold. ‘w/o rewriting’ and ‘w/o retrieval’ refer to results without the rewriting step and retrieval step in SSE-CoT, respectively. ‘ w/o both’ refers to the exclusion of both steps
SSE-CoT Ablation. Ablation studies were performed on six benchmark datasets to assess the influence of the components of the SSE-CoT framework. We developed two variations of the QLFQ as follows.
  • SSE-CoT+w/o rewriting: Omission of the rewriting step, using the concatenation of the original input with the retrieved text as input.
  • SSE-CoT+w/o retrieval: Omission of the retrieval step, employing rewritten original inputs as input.
  • SSE-CoT+w/o both: Utilizing the original input without modification.
@@
4.7 Analysis of In-Context Learning for SSE-CoT
In addition to the supervised fine-tuning (SFT) paradigm, in-context learning has gained popularity. Four LLMs, ChatGLM, LLaMA2-7B, LLaMA2-13B, and GPT-3 were selected for experimentation under the zero-shot and one-shot settings. Given the rigorous demands of in-context learning, the SSE-CoT was tested solely on the GPT-3. We manually selected a sample from each category in a one-shot setting and constructed the SSE-CoT as the context. Specific examples of the input formats are provided in Appendix B. The experimental results are illustrated in Fig 5.
@@ 4.7 Analysis of In-Context Learning for SSE-CoT
Zero-shot setting. In the zero-shot setting analysis conducted across three datasets, the in-context learning abilities of the four models were ranked as follows: GPT-3 outperformed LLaMA2-13B, which, in turn, surpassed both ChatGLM and LLaMA2-7B. This hierarchy can be attributable to in-context learning capabilities inherent in larger models, with increased parameters correlating with enhanced performance. Applying our SSE-CoT method to GPT-3 resulted in performance gains across the board. For instance, the ACC on the Ohsumed dataset increased from 51.6% to 52.9%. These findings indicate that our approach is not limited to the SFT but is also applicable to the context learning paradigm.
@@
Comparing the results of the zero-shot and one-shot tasks, contrary to our initial expectations, the hypothesis that providing task-relevant examples enhances the comprehension and task response of LLMs was not supported. The findings indicate that only GPT improves performance in one-shot learning compared with zero-shot learning on two of the datasets. In contrast, a decline is observed in the remaining datasets. Further analysis of the model outputs suggested that LLM with approximately 10 billion parameters tended to demonstrate a reduced capacity for processing instructions with greater input complexity. In contrast, the GPT consistently exhibits a robust understanding of instructions and sustains its performance despite increased context length.
Comparison of the results of SFT and in-context learning paradigm. GPT-3, with over a hundred billion parameters, exhibits a remarkable in-context learning performance. In the one-shot setting, GPT-3 exceeded CDMT, and GPT-3+SSE-CoT nearly matched the supervised SSE-CoT method. However, for models with billions of parameters, the SFT paradigm remained predominant.
4.8 Analysis of different base models for CDMT
In our CDMT method, the initial stage involves the extraction of rationales with the assistance of an LLM, followed by a fine-tuning phase that requires a smaller model. Therefore, we investigated the outcomes of employing the same LLM with various smaller models and the results of utilizing the same smaller model with different LLMs.
@@ 4.8 Analysis of different base models for CDMT
Different smaller models. As shown in Fig 6, we utilized LLaMA2-13B as the LLM and three versions of Flan-T5-Base, Flan-T5-Large, and Flan-T5-XL as smaller models, with parameter counts of 250M, 780M, and 3 B, respectively. It is apparent that, with an increase in the parameter size of the smaller models, there is a corresponding enhancement in both the ACC and F1 metrics, which suggests a correlation between the capacity of the smaller model and its performance. However, this relationship is not strictly linear. For instance, in the Ohsumed dataset, the increase in ACC from the 250M model to the 780M model does not reflect the expected proportional improvement, indicating diminishing returns as the model size increases.
@@
4.9 Analysis of different prompts for CDMT
@@
Table 6: Evaluation of Different Prompts in CDMT: the efficacy of ECCA demonstrated through significant outcomes.
Table 6 highlights several interesting trends. Prompts enhanced performance across all measured metrics relative to their absence. Furthermore, our method outperformed the less semantically rich prompt1 and demonstrated greater accuracy than prompt2 in the two datasets. Although our approach does not outperform prompt2 on the MR dataset in terms of ACC, it offers considerable benefits. It obviates the necessity for creating manual, dataset-specific prompts, which yield notably shorter input lengths than prompt2 by simply appending label names, benefiting processing efficiency and model scalability.
@@ 4.10 Analysis of training size
The results presented in Fig 7 show that the proposed method significantly outperforms the other two models in low-resource environments. The data reveal a consistent increase in test accuracy correlated with the rising proportion of training data for both the MR and TagMyNews datasets, consistent with supervised learning principles. Additionally, the acceleration or plateau in ACC improvement suggests diminishing returns in the model performance beyond a certain data threshold. Notably, the ACC of SHINE and ChatGLM approach parity, whereas our methods consistently outperform them, indicating that, with adequate data, the inherent strengths of the model become increasingly influential.
4.11 Analysis of time complexity
The results presented in Table 7 assess the time complexity of four methods using the TagMyNews dataset, which includes a training set of 140 samples and a test set of 24,600 samples. KG denotes constructs a knowledge graph and RG stands for rationale generation. As shown, SHINE constructs a knowledge graph in 340 seconds, whereas CDMT generates rationales using a large language model in only 13 seconds. In the training phase, SHINE completes its process in 94 seconds, while SSE-CoT requires significantly more time at 1859 seconds. For inference time, CDMT and SSE-CoT require 326 seconds and 804 seconds, respectively.
Traditional methods such as SHINE excel in training speed compared to LLM-based approaches. However, in practical applications, inference time is crucial. SHINE requires reconstruct and retrain each new test sample, making its inference time equal to its construction and training times combined. In contrast, CDMT’s pipeline design is more efficient, offering considerable advantages. While SSE-CoT inference is slower than SHINE, it remains feasible for large-scale deployment.
RG/KG(s) Train(s) Inference(s)
SHINE 340 94 434
ChatGLM 1446 752
CDMT 13 577 326
SSE-CoT 1859 804
Table 7: Evaluation of time complexity on TagMyNews Dataset.
@@ 5 Conclusion
In this study, we developed and evaluated novel methods to improve Short Text Classification (STC) using Large Language Models (LLMs) and a Chain-of-Thought (CoT) processing approach. The Semantic and Syntactic Enrichment CoT (SSE-CoT) method breaks down the STC tasks into four steps, facilitating thorough comprehension and management of short texts. This approach outperforms traditional models by providing a level of semantic and syntactic analysis that was not achievable with Graph Convolutional Networks (GCNs). In parallel, acknowledging the challenges faced in resource-constrained sectors like finance and healthcare, we introduced the CoT-Driven Multi-Task learning (CDMT) framework. This method leverages insights from LLMs and adapts them for smaller models, improving their efficiency and effectiveness through targeted fine-tuning and multi-task learning strategies. Comprehensive experiments were conducted across six prevalent datasets to evaluate the effectiveness of the proposed methods. Experimental results indicate that the proposed methods significantly outperformed established baselines. However, this complex task remains partially unresolved because its performance on domain-specific datasets remains suboptimal. Future research should prioritize the integration of LLMs with additional knowledge sources to further refine the proposed methodologies.
@@ 6 Limitations
Our study validated the effectiveness of the proposed method through carefully designed experiments, although several limitations were encountered. Constraints related to hardware resources and time restricted the number of LLMs employed in this study. Moreover, the rapid advancements in LLMs technologies pose challenges in maintaining up-to-date comparative analyses. A significant limitation is the increased time complexity during the training and inference phases of our methods, a typical trade-off when using advanced models. Despite these challenges, we made considerable efforts to ensure the robustness and relevance of our findings. In future work, we aim to expand our research by incorporating more diverse LLMs and enhancing the efficiency of our algorithms. This approach seeks to address the current limitations related to hardware and time constraints while enhancing the scalability of our methods. Additionally, we intend to update our comparative analyses regularly to align with the rapid advancements in LLMs technologies.
@@ Appendix B In-Context Learning input apllications
Appendix C Different Prompts for CDMT
{mdframed}
[ frametitle=Prompt1, frametitlerule=true, frametitlebackgroundcolor=gray!60, linewidth=1.5pt, roundcorner=5pt, backgroundcolor=gray!10, linecolor=black, align=center, userdefinedwidth=0.45] Categorize this text: ‘wal-mart buys social media firm kosmix’.
{mdframed}
[ frametitle=Prompt2, frametitlerule=true, frametitlebackgroundcolor=gray!60, linewidth=1.5pt, roundcorner=5pt, backgroundcolor=gray!10, linecolor=black, align=center, userdefinedwidth=0.45] Given the short text ‘wal-mart buys social media firm kosmix’, classify it into one of the categories. The categories are health, sport, entertainment, business, sci_tech, U.S. and world.
Appendix D Case study
Figure 12 presents a case study where our CDMT framework optimized a smaller model to enhance short text classification. Initially, the model incorrectly classified the phrase ‘Del Potro says make French Open’ as ‘world’ due to poor semantic and syntactic comprehension. Following optimization, the model’s accuracy improved markedly, correctly categorizing the phrase under ‘sports’. This enhancement illustrates the impact of advancing the model’s understanding of semantics and syntax on classification accuracy and reliability.
Figure 12: case study of CDMT